\begin{thebibliography}{9}

%\bibitem{AkPe} S. Akiyama and A. Peth{\H o},
%`On the distribution of polynomials with bounded roots II. Polynomials
%with integer coefficients', {\text Preprint\/}, 2012.
%%
%%










%%


\bibitem{ABS} A. G. Abercrombie, W. D. Banks and I.~E.~Shparlinski,
`Arithmetic functions on Beatty sequences',
\textit{Acta Arith.\/}, {\bf 136} (2009), 81--89.

\bibitem{Akb} Y. Akbal,
`Friable values of Piatetski-Shapiro sequences',
\textit{Proc. Amer. Math. Soc.\/}, {\bf 145} (2017),   4255--4268.

 \bibitem{BaBa}
R.~C.~Baker and W.~D.~Banks,
`Character sums with Piatetski--Shapiro sequences',
\textit{Quart. J.  Math.}, {\bf  66} (2015), 393--416.

\bibitem{BBBSW}
R.~C.~Baker, W.~D.~Banks, J.~Br\"udern, I.~E.~Shparlinski and A.~Weingartner,
`Piatetski--Shapiro sequences',
\textit{Acta Arith.}, {\bf 157} (2013),  37--68.

\bibitem{BBGY}
R.~C.~Baker, W.~D.~Banks, V.~Z.~Guo and A. M. Yeager,
`Piatetski--Shapiro primes from almost primes',
\textit{Monatsh Math.\/}, {\bf  174} (2014), 357--370.


 \bibitem{BaLi}
R.~C.~Baker and L. Zhao,  `Gaps between primes in Beatty sequences',
\textit{Acta Arith.}, {\bf 172} (2016),   207--242.

\bibitem{BGS}
W.~D.~Banks,  V.~Z.~Guo and  I.~E.~Shparlinski,
`Almost primes of the form $\fl{p^c}$'',
\textit{Indag. Math.\/}, {\bf 27}  (2016),  423--436.

\bibitem{Bor}
O.~Bordell\`es, L.~Dai, R.~Heyman, H.~Pan and I.~E.~Shparlinski,
`On a sum involving the Euler function',
{\text J. Number Theory\/}, \textbf{202} (2019),  278--297.

\bibitem{Bor2}
O.~Bordell\`es,
`On certain sums of number theory',
\textit{Preprint}, 2020, available at arXiv:2009.05751 [math.NT].

\bibitem{Bor3}
O. Bordell\`{e}s,  \textit{Arithmetic tales (2nd edition)},
Springer, 2020.


%\bibitem{Bord2}  O. Bordell\`{e}s,
%`Short interval results for certain arithmetic functions',
%\textit{Int. J. Number Theory},  {\bf 14} (2018), 535--548.
%
%\bibitem{Bour}  J. Bourgain, `Decoupling, exponential sums
%and the Riemann zeta function', \textit{J. Amer. Math. Soc.},
%{\bf 30} (2017), 205--224.
%
%\bibitem{BouWat}  J. Bourgain and N. Watt,
%`Mean square of zeta function, circle problem and divisor problem
%revisited', \textit{Preprint}, 2017, available at arXiv:1709.04340 [math.NT].
%
%\bibitem{DDM} J.-M. Deshouillers,  M. Drmota,  and  J. F. Morgenbesser,
%`Subsequences of automatic sequences indexed by $\fl{n^c}$ and correlations',
%\textit{J. Number Theory}, {\bf 132} (2012),  1837--1866.

\bibitem{Che}
S.~Chern,
`Note on sums involving the Euler function' (English summary),
\textit{Bull. Aust. Math. Soc.}, {\bf 100} (2019), 194--200.

\bibitem{GuNe} A.~M.~G{\"u}lo{\v g}lu and C.~W.~Nevans,
`Sums with multiplicative functions over a  Beatty sequences',
\textit{Bull.  Austral. Math.  Soc.}, {\bf 78} (2008), 327--334.



\bibitem{Harm} G. Harman,  `Primes in Beatty sequences in short intervals',
\textit{Mathematika}, {\bf 62} (2016), 572--586.

\bibitem {Hey}
R.~Heyman,
`Cardinality of a floor function set',
{\text Integers\/}, \textbf{19} (2019), A67.

\bibitem{Liu}
K.~Liu, J.~Wu and Z.~Yang,
`On some sums involving the integral part function',
\textit{Preprint}, 2021, Available at arXiv:2109.01382 [math.NT]

\bibitem{Liu2}
K.~Liu, J.~Wu and Z.~Yang,
`A variant of the Prime Number Theorem',
\textit{Preprint}, 2021, Available at arXiv:2105.10844 [math.NT]






\bibitem{LSZ}
K. Liu,  I. E. Shparlinski and T. Zhang,
`Squares in Piatetski--Shapiro sequences',
\textit{Acta Arith.}, {\bf 181} (2017),  239--252.

\bibitem{Ma}
J.~Ma and J.~Wu,
`On a sum involving the Mangoldt function' (English summary),
\textit{Period. Math. Hungar.}, {\bf 83} (2021), 39--48.


\bibitem{Ma2}
J.~Ma and H.~Sun,
`On a sum involving certain arithmetic functions and the integral part function',
\textit{Preprint}, 2021, available at arXiv:2109.02924 [math.NT]

\bibitem{Morg}    J. F. Morgenbesser,
`The sum of digits of $\fl{n^c}$',
\textit{Acta Arith.}, {\bf 148} (2011),  367--393.


%\bibitem{ramak} O. Ramar\'{e} and P. Akhilesh, `Explicit averages of non-negative multiplicative functions: going beyond the main term', \textit{Coll. Math.}, {\bf 147} (2017), 1--39.

\bibitem{Ros}
J.~Rosser, J.~Barkley and L.~Schoenfeld,
`Approximate formulas for some functions of prime numbers',
\textit{Illinois J. Math.}, {\bf 6} (1962) 64--94.

%\bibitem{Shu}
%A.~V.~Shubin,
%`Fractional parts of the function $x/n$',
%\textit{Math. Notes}, {\bf 100} (2016), 774--756.

%   \bibitem{sur} D. Suryanarayana and  P. Subrahmanyam, The maximal $k$-full divisor of an integer, \textit{Indian J. Pure Appl. Math.}, {\bf 12} (1981), 175--190.
\bibitem{Stu}
J.~Stucky,
`The fractional sum of small arithmetic functions',
\textit{Preprint}, 2021, available at arXiv:2106.14142 [Math.NT].








%\bibitem{Ten} G.~Tenenbaum,  \textit{Introduction to analytic and
%probabilistic number theory}, Grad. Studies Math., vol.~163, Amer. Math. Soc., 2015.
%
%\bibitem{tre14} E. Trevi\~{n}o, The least $k$-th power non-residue, \textit{J. Number Theory}, {\bf 149} (2015), 201--224.
%
%\bibitem{tre15} E. Trevi\~{n}o, The Burgess inequality and the least $k$-th power non-residue, \textit{Int. J. Number Theory}, {\bf 11} (2015), 1653--1678.
%
%\bibitem{Vaal} J. D. Vaaler, `Some extremal functions in Fourier analysis',
%\textit{Bull. Amer. Math. Soc.}, {\bf 12} (1985), 183--215.
%
%
%\bibitem{Wal}
%A.~Walfisz,
%\emph{Weylsche Exponentialsummen in der neueren Zahlentheorie},
%Berlin (1963)

\bibitem{Wu}
J.~Wu,
`On a sum involving the Euler totient function',
\textit{Indag. Math. (N.S.)}, {\bf 30} (2019), 536–541.

\bibitem{Zha}
W.~Zhai,
`On a sum involving the Euler function',
\textit{J. Number Theory}, {\bf 211} (2020), 199--219.

\bibitem{Zha2}
W.~Zhai,
`Corrigendum to "On a sum involving the Euler function''',
\textit{J. Number Theory}, {\bf 222} (2021), 423--425.

\bibitem{Zhao}
F.~Zhao and J.~Wu,
`On a Sum Involving the Sum-of-Divisors Function',
\textit{Journal of Mathematics}, {\bf 2021}, 2021.
%
%\bibitem {BaHa2} R.~C.~Baker and G.~Harman,
%`Small remainder of a vector to suitable modulus', {\text Math. Zeit.\/},
%\textbf{221} (1996), 59--71.
%%
%\bibitem {Bak} R.~C.~Baker,~G.~Harman and J.~Pintz,
%`The difference between consecutive primes, II', {\text Proc. Lond. Math. Soc.\/},
%\textbf {83} (2001), 532--562.
%%
%\bibitem{Bal}
%U.~Balakrishnan and Y.~F.~S. P$\acute{\textrm{e}}$termann,
%`The Dirichlet series of $\zeta(s)\zeta^\alpha(s+1)f(s+1)$: On an error term associated with its coefficients',
%{\text Acta Arith.\/}, \textbf{75}~(1996), 39--69.
%%
%\bibitem{Bis} A.~Bishnoi and S.~K.~Khanduja,
%`On Eisenstein-Dumas and generalized Sch$\ddot{\textrm{o}}$nemann polynomials',
%{\text Comm. Algebra\/}, \textbf{38} (2010), 3163--3173.
%%
%\bibitem{Bon} N.~C.~Bonciocat,
%`Sch$\ddot{\textrm{o}}$nemann-Eisenstein-Dumas-type irreducibility conditions that use arbitrarily many prime numbers',
%Preprint, 2014, available at arXiv:1304.0874 [math.NT].
%%
%\bibitem{BMGVBO}
%F. E. Brochero Mart\'inez, C. R. Giraldo Vergara and L. Batista de Oliveira,
%`Explicit factorization of $x^n-1\in \F_q[x]$',
%Preprint, 2014, available at arXiv:1404.6281 [math.NT].
%%
%\bibitem{ChNg} Y. Chen and P. Q. Nguyen,
%`Faster algorithms for approximate common divisors:
%Breaking fully-homomorphic-encryption challenges over the integers',
%{\text Advances in Cryptology-EUROCRYPT 2012\/}, Lecture Notes Comput. Sci., \textbf{7237},
%Springer, Berlin, 2012,  502--519.
%%
%\bibitem{Cho}
%S.~D.~Chowla,
%`An order result involving the Euler $\varphi$-function',
%{\text J. Ind. Math. Soc. \textrm{Old Ser.}\/}, \textbf{18}~(1929--1930), 138--141.
%%
%\bibitem{Chu}
%K.~Chung, Y.~Kalai and S.~Vadhan,
%`Improved delegation of computation using fully homomorphic encryption',
%{\text Advances in Cryptography-CYPTO 2010\/}, Lecture Notes Comput. Sci., \textbf{6223}, Springer, Berlin, (2010), 483--501.
%%
%\bibitem{Coh}
%S.~D.~Cohen,
%`The distribution of the Galois groups of integral polynomials',
%{\text Illinois J. of Math.}, \textbf{23}  (1979), 135--152.
%%
%\bibitem{CoHe}
%H. Cohn and N.  Heninger, `Approximate common divisors via lattices',
%Preprint, 2011,
%available at arXiv:1108.2714 [math.NT].
%%
%\bibitem{Cohen} S. D. Cohen,
%`Explicit theorems on generator polynomials',
%{\text Finite Fields Appl.\/}, \textbf{11} (2005),   337--357.
%%
%\bibitem{Coo} G.~Cooperman, S.~Feisel, J.~von zur Gathen and G.~Havas,
%`GCD of many integers (extended abstract)',
%{\text Computing and combinatorics 1999\/},
%Lecture Notes Comput. Sci., \textbf{1627}, Springer, Berlin, 1999, 310--317.
%%
%\bibitem{Cor} J.~Coron and A.~May,
%`Deterministic polynomial-time equivalence of computing the RSA secret key and factoring',
%{\text J. Cryptology\/}, \textbf{20} (2007),   39--50.
%%
%\bibitem{Cox}
%D.~A.~Cox,
%`Why Eisenstein proved the Eisenstein criterion and why Sch$\ddot{\textrm{o}}$nemann discovered it first',
%{\text Amer. Math. Monthly}, \textbf{118} (2011), 3--21.
%%
%\bibitem{Diet} R.~Dietmann,
%`On the distribution of Galois groups',
%{\text Mathematika\/}, \textbf{58} (2012),  35--44.
%%
%\bibitem{vDGHV} M.~van Dijk, C.~Gentry, S.~Halevi and V.~Vaikuntanathan,
%`Fully homomorphic encryption over the integers',
%{\text Advances in cryptology-EUROCRYPT 2010\/}, Lecture Notes Comput. Sci., \textbf{6110}, Springer, Berlin, 2010, 24--43.
%%
%\bibitem{DinPeiSal} C.~Ding, D.~Pei and A.~Salomaa, \emph{The Chinese remainder theorem}, World Scientific, 1996.
%%
%\bibitem{DoJo} D.~E.~Dobbs and L.~E.~Johnson,
%`On the probability that Eisenstein's criterion applies to an arbitrary irreducible polynomial',
%{\text 3rd Intern. Conf. Advances in Commutative Ring Theory\/}, Lecture Notes in Pure and Appl. Math., \textbf{205}, Dekker, New York, 1999, 241--256.
%%
%\bibitem{Dot} E.~Dotti and G.~Micheli,
%`Eisenstein polynomials over function fields',
%Preprint, 2015, avaiable at arXiv:1506.05380 [math.NT].
%%
%\bibitem{DIT} F. Dress, H.~Iwaniec, and G. Tenenbaum,
%`Sur une somme li{\'e}e {\`a} la fonction de M{\"o}bius',
%{\text J. Reine Angew. Math.\/}, \textbf {340} (1983), 53--58.
%%
%\bibitem{Dub}
%A.~Dubickas,
%`Polynomials irreducible by Eisenstein's criterion',
%{\text Appl. Algebra Engin. Comm. Comput.}, \textbf{14} (2003), 127--132.
%%
%\bibitem{Dum}
%G.~Dumas,
%`Sur quelques cas d'irr$\acute{\textrm{e}}$ductibilit$\acute{\textrm{e}}$ des polynomes $\grave{\textrm{a}}$ coefficients rationnels',
%{\text J. Math. Pures Appl.\/}~{\textbf(6)\/}, \textbf{2}~(1906), 191--258.
%%
%\bibitem{Eis}
%F.~G.~Eisenstein, \emph{Mathematische werke, Vol. II}, AMS Chelsea,
%Providence, RI, 1989, 536--555.
%%
%\bibitem{EGB} M. Elkadi, A. Galligo and  T. L. Ba,
%`Approximate GCD of several univariate polynomials with small degree perturbations',
%{\text J. Symbolic Comput.\/}, \textbf{47} (2012),  410--421.
%%
%\bibitem{Erd} P. Erd{\H o}s,
%`On an elementary problem in number theory',
%{\text Canadian Math. Bull.\/}, \textbf{1}  (1958), 5--8.
%%
%\bibitem{Fer} J.~L.~Fern$\acute{\textrm{a}}$ndez and P.~Fern$\acute{\textrm{a}}$ndez,
%`Equidistribution and coprimality',
%Preprint, 2013, avaiable at arXiv:1310.3802 [math.NT].
%%
%\bibitem{Fre}
%T.~Freiberg,
%`The probability that 3 positive integers are pairwise noprime',
%(unpublished manuscript).
%%
%\bibitem{Gat} J.~von zur Gathen and J\"{u}rgen Gerhard,
%\emph{Modern computer algebra, (3rd edition)\/}, Cambridge University Press, 2013.
%%
%\bibitem{vzGMS} J. von zur Gathen, M. Mignotte and  I. E. Shparlinski,
%`Approximate polynomial GCD: Small degree and small height perturbations',
%{\text J. Symbolic Comput.\/}, \textbf{45} (2010),   879--886.
%%
%\bibitem{vzGSh} J. von zur Gathen and I. E. Shparlinski,
%`GCD of random linear combinations',
%{\text  Algorithmica\/}, {\textbf {46}} (2006), 137--148.
%%
%\bibitem{Gau} C.~F.~Gauss,
%\emph{Disquisitiones arithmeticae (English Edition)}, Springer-Verlag,
%New York, 1986.
%%
%\bibitem{Gen} C.~Gentry,
%`Fully homomorphic encryption using ideal lattices',
%{\text Proc. of the ACM Intern. Symp. on Theory of Comp. 2009\/},
%ACM, New York, 2009, 169--178.
%%
%\bibitem{GrigTen} D.  Grigoriev  and G. Tenenbaum,
%`A low complexity probabilistic test for integer multiplication',
%{\text J.  Complexity\/}, \textbf{26} (2010), 263--267.
%%
%\bibitem{Got} J.~Goth,
%`Minimizing Non-interactive Zero-Knowledge Proofs Using Fully
%Homomorphic Encryption', {\text IACR Cyptology\/}, ePrint Archive 2011, (2011).
%%
%\bibitem{Gro} T.~H.~Gronwall,
%`Some asymptotic expressions in the theory of numbers',
%{\text Trans. Amer. Math. Soc.\/}, \textbf{14}~(1913), 113--122.
%%
%\bibitem{Har} G.~H.~Hardy and E.~M.~Wright,
%\emph{An introduction to the theory of numbers (6th Edition)}, Oxford University Press,
%Oxford, 2008.
%%
%\bibitem{Hara} F.~Harary and E.~M.~Palmer,
%\emph{Graphical Enumeration}, Academic Press,
%New York, 1973.
%%
%%
%\bibitem{Hey}
%R.~Heyman,
%`Pairwise non-coprimality of triples',
%Preprint, 2014, available at arXiv: 1309.5578 [math.NT].
%%
%\bibitem{HeSh}
%R.~Heyman and I.~E.~Shparlinski,
%`On the number of Eisenstein polynomials of bounded height',
%{\text Appl. Algebra Engrg. Comm. Comput.\/}, \textbf{24} ~(2013), 149--156.
%%
%\bibitem{Hey2}
%R.~Heyman and I.~E.~Shparlinski,
%`On shifted Eisenstein polynomials',
%{\text Period. Math. Hungar.\/}, \textbf{69}~(2014), 170--181.
%%
%\bibitem{Hey3}
%R.~Heyman,
%`On the number of polynomials of bounded height that satisfy the Dumas criterion',
%{\text J. Integer Seq.\/}, \textbf{17} ~(2014), 14.2.4.
%%
%\bibitem{Hey4}
%R.~Heyman and I.~E.~Shparlinski,
%`On the GCD of shifted sets',
%{\text J. Number Theory\/} \textbf{154}~(2015), 63--73.
%%
%\bibitem{Hey5}
%R.~Heyman and I.~E.~Shparlinski,
%`Counting irreducible binomials over finite fields',
%Preprint, 2105, avaiable at arXiv: 1504.01172 [math.NT].
%%
%\bibitem{How} N.~Howgrave-Graham,
%`Approximate integer common divisors',
%\text{Cryptology and lattices 2001\/},
%Lecture notes in Comput. Sci.,
%\textbf{2146}, Springer, Berlin, 2001, 51--66.
%%
%\bibitem{Hu1}
%J.~Hu,
%`The probability that random positive integers are $k$-wise relatively prime',
%{\text In. J. Number Theory\/}, \textbf{9}~(2013), 1263--1271.
%%
%\bibitem{Hu2}
%J.~Hu,
%`Pairwise relative primality of positive integers',
%Preprint, 2014, available  at arXiv:1406.3113 [math.NT].
%%
%\bibitem{HuPl}
%C. Huck and  P. A. B. Pleasants, `Entropy and diffraction of the
%$k$-free points in $n$-dimensional lattices',
%{\text Discrete Comput. Geom.\/}, \textbf{50} (2013), 39--68.
%%
%\bibitem{Hucz} S. Huczynska,
%`Existence results for finite field polynomials with specified properties',
%\emph{Finite fields and their applications: character sums and polynomials},
%De Gruyter, Berlin, 2013, 65--87.
%%
%\bibitem{Iw} H.~Iwaniec,
%`On the problem of Jacobsthal', {\text
%Demonstratio Math.\/}, \textbf{11} (1978), 225--231.
%%
%\bibitem{IwKow} H. Iwaniec and E. Kowalski,
%\emph{Analytic number theory}, Amer.  Math.  Soc.,
%Providence, RI, 2004.
%%
%\bibitem{Jur} M.~Jur$\acute{\textrm{a}}$$\check{\textrm{s}}$,
%`Eisenstein-Dumas criterion for irreducibility of polynomials and projective transformations of the independent variable',
%{\text JP Jour. Algebra, Number Theory \& Appl.\/}, \textbf{6}, 2006, 221--236.
%%
%\bibitem{Kac}
%M.~Kac,
%Statistical independence in probability, analysis and number theory,
%{\text The Carus Mathematical Monographs, No. 12.\/} The Math. Assoc. of America, 1959.
%%
%\bibitem{Kat} V.~J.~Katz, \emph{A history of mathematics (brief edition)}, Pearson/Addison-Wesley, 2003.
%%
%\bibitem{Knu} D.~E.~Knuth,
%\emph{The Art of computer programming, Vol 2. Seminumerical algorithms (3rd Edition)},
%Addison Wesley, Boston, 1998.
%%
%\bibitem{LLL} A.~K.~Lenstra, H.~W.~Lenstra and L.~Lov{\'a}sz,
%`Factoring polynomials with rational coefficients',
%{\text Mathematische Annalen\/}, \textbf{261} (1982), 515--534.
%%
%\bibitem{LiNi} R.~Lidl and H.~Niederreiter,
%\emph{Finite fields}, Addison-Wesley 1983.
%%
%\bibitem{Liu}
%J.~Y.~Liu,
%`On an error term of Chowla, I',
%{\text J. Number Theory\/}, \textbf{64}~(1997), 20--35.
%%
%\bibitem{Lop}
%A.L$\grave{\textrm{o}}$pez, E.~Tromer and V.~Vaikuntanathan,
%`On-the-fly multiparty computation on the cloud via multikey fully homomorphic encryption'
%{\text Proc. of the 44th annual ACM symposium on theory of computing\/}, ACM, (2012), 1219--1234.
%%





%\bibitem{Mahler}
%K.~Mahler,
%`An inequality for the discriminant of a polynomial',
%{\text Michigan Math. J.\/}, \textbf{11} (1964), 257--262.
%%
%\bibitem{MaZi} A. Masuda and M. E. Zieve,
%`Permutation binomials over finite fields',
%{\text Trans. Amer. Math. Soc.\/}, \textbf{361} (2009), 4169--4180.
%%
%\bibitem{Maz}
%G.~Maze, J.~Rosenthal and U.~Wagner,
%`Natural density of rectangular unimodular integer matrices',
%{\text In. Linear Algebra Appl.\/}, \textbf{434}~(2011), 1319--1324.
%%
%\bibitem{Mic} D.~Micciancio,
%`A first glimpse of cryptography's holy grail', {\text
%Comm. of the Assoc. of Comp. Machinery\/}, \textbf{53} (2010), 96--96.
%%
%\bibitem{Mic2}
%G.~Micheli and R.~Schnyder,
%`The density of shifted and  affine Eisenstein polynomials',
%Preprint, 2015, available at arXiv:1507.02753 [math.NT].
%%
%\bibitem{Mik}
%H.~Mikawa, `On primes in arithmetic progressions', {\text Tsukuba
%J.\ Math.\/}, \textbf{25} (2001), 121--153.
%%
%\bibitem{Mor}
%P.~Moree,
%`Counting carefree couples',
%Preprint, 2015, available at arXiv:0510003 [math.NT].
%%
%%\bibitem{Mori} K. Morita, {\textit Applied fourier transform \/}, Ohmsha, 1995.\comm{nowhere?}
%%
%\bibitem{Na} K. Nagasaka, `Approximate polynomial GCD over integers',
%{\text J. Symbolic Comput.\/}, \textbf{46} (2011),  1306--1317.
%%
%\bibitem{Nym} J.~E.~Nymann,
%`On the probability that $k$ positive integers are relatively prime',
%{\text J. Number Theory\/}, \textbf{4}~(1972), 469--473.
%%
%\bibitem{Osa} H.~Osada,
%`The Galois groups of the polynomials $x^n+ax^l+b$',
%{\text  J. Number Theory\/}, \textbf{25} (1987), 230--238.
%%
%\bibitem{Poll} P.~Pollack,  `Irreducible polynomials with several prescribed coefficients',
%{\text Finite Fields Appl.\/}, \textbf{22} (2013), 70--78.
%%
%\bibitem{Poon}
%B.~Poonen, `Squarefree values of multivariable polynomials',
%{\text  Duke Math. J.\/}, \textbf{118} (2003), 353--373.
%%
%\bibitem{Pra} V.~V.~Prasolov, {\textit{Polynomials}\/}, Springer, 2004.
%%
%%\bibitem{Rud} P.~S. Rudman,
%%\emph{How mathematics happened: The first 50,000 years}, Prometheus Books, 2007.
%%
%\bibitem{SaMa} S. Sarkar and S. Maitra, `Approximate integer common
%divisor problem relates to implicit factorization',
%{\text IEEE Trans. Inform. Theory\/}, \textbf{57} (2011), 4002--4013.
%%
%\bibitem{Shoup} V. Shoup,
%`New algorithms for finding irreducible polynomials over finite fields',
%{\text Math. Comp.\/}, \textbf{54} (1990), 435--447.
%%
%\bibitem{Sur} D.~Suryanarayana,
%`Asymptotic formula for $\sum_{n \le x} \frac{\mu^2(n)}{n}$',
%{\text Indian J. Math.\/}, \textbf{9} (1967), 543--545.
%%
%\bibitem{Ten} G.~Tenenbaum,
%\emph{Introduction to analytic and probabilistic number theory\/}, Cambridge University Press, 1995.
%%
%\bibitem{Tot} L.~T$\acute{\textrm{o}}$th,
%`The probability that $k$ positive integers are pairwise relatively prime',
%{\text Fibonacci Quart.\/}, \textbf{40}~(2002), 13--18.
%%
%\bibitem{Vau}
%R. C. Vaughan, `An elementary method in prime number theory',
%{\text Acta Arith.\/}, \textbf{37} (1980), 111--115.
%%
%\bibitem{Wre} J.~W.~Wrench Jr.,
%`Evaluation of Artin's constant and the twin prime conjecture',
%{\text Math Comp.\/}, \textbf{15}~(1961), 396--398.
%%
%\bibitem{Zyw} D. Zywina, `Hilbert's irreducibility theorem and the larger sieve',
%{\text  Preprint\/}, 2010, available from arXiv/1011.6465 [math.NT].



\end{thebibliography}

\RequirePackage[l2tabu, orthodox]{nag}
\documentclass[10pt, draft]{article}

\usepackage{amsmath,amssymb,amsbsy,amsfonts,amsthm,latexsym,amsopn,amstext,
                            amsxtra,euscript,amscd}
\usepackage{booktabs}
\usepackage{enumitem}
\usepackage{tikz}
\usepackage[norefs,nocites]{refcheck}
\newtheorem{thm}{Theorem}
\newtheorem{lem}{Lemma}
\newtheorem{lemma}[thm]{Lemma}

\newtheorem{conj}[thm]{Conjecture}

\def\\{\cr}
\def\({\left(}
\def\){\right)}
\def\[{\left[}
\def\]{\right]}

\def\<{\langle}
\def\>{\rangle}
\def\fl#1{\left\lfloor#1\right\rfloor}

\def\cC{\mathcal C}
\def\cF{\mathcal F}

\def\cG{\mathcal G}

\def\notdivides{\mathrel{\kern-3pt\not\!\kern3.5pt\bigm|}}

\begin{document}


\title{\Large \textbf{Primes in floor function sets}}
\author{
\scshape {RANDELL HEYMAN}  \\
School of Mathematics and Statistics,\\ University of New South Wales \\
Sydney, Australia\\
\texttt {randell@unsw.edu.au}
}

\maketitle


\begin{abstract}
Let $x$ be a positive integer. We give an asymptotic formula for the number of primes in the set $\{\fl{x/n}, 1 \le n \le x\}$ and give some related results.
\end{abstract}


\section{Introduction}
There is an extensive body of research on
arithmetic functions with integer parts of real-valued functions, most commonly, with
 Beatty $\fl{\alpha n + \beta}$  sequences, see, for
 example,~\cite{ABS,BaBa,BaLi,GuNe,Harm}, and
 Piatetski--Shapiro $\fl{n^\gamma}$ sequences,
 see, for example,~\cite{Akb,BBBSW,BBGY,BGS,LSZ,Morg},
 with real $\alpha$, $\beta$ and $\gamma$.

 Recently there has been much research on sums of the form
 \begin{align}
 \label{eq:f sum}
 \sum_{n \le x}f\(\fl{\frac{x}{n}}\),
 \end{align}
 where throughout $x$ is a positive integer, $f$ is an arithmetic function and $\fl{\cdot}$ is the floor function.
 In \cite{Bor} the authors used exponential sums to find asymptotic bounds and formulas for various classes of arithmetic functions. Subsequent papers by various authors have mainly been focussed on improvements in exponential sums techniques (see \cite{Bor2,Che, Liu,Ma,Ma2,Stu,Wu,Zha,Zha2,Zhao}).

It is natural to examine more fundamental questions about the set$\{\fl{x/n}: 1 \le n \le x\}$. In \cite{Hey} an exact formula  for the cardinality of this set was given.
In this paper we count primes in this floor function set. Let
 $$\cG(x)=\left\{\fl{\frac{x}{n}}: 1 \le n \le x, \fl{\frac{x}{n}}\text{ is prime}\right\},$$
 and in particular $G(x):=|\cG(x)|$ This can estimated using exponential sums as follows:
\begin{thm}
 \label{thm:G(x)}
 We have
 $$G(x)=\frac{4 \sqrt{x}}{\log x}+O\(\frac{ \sqrt{x}}{(\log x)^2}\).$$
 \end{thm}

 The OEIS sequence A068050 attributes to Adams-Watters the statement that for $p$ prime not equal to 3 we have $G(p)=G(p-1)+1$. A proof does not seem evident. We prove the following:
\begin{thm}
\label{thm:adams}
Let $x$ be any prime not equal to 3. Then $G(x)=G(x-1)+1$.
\end{thm}
It is possible to link up $G(x)$ and $G(x-1)$ for some other classes of $x$ as follows:
 \begin{thm}
 \label{thm:G(x) semiprime}
Let $x=pq$ with $p,q$ odd primes, not necessarily distinct.
 Then
 $$G(x)=G(x-1)+1.$$
 \end{thm}
 These relationships between $G(x)$ and $G(x-1)$ may generalise, but with considerable difficulties.  For example, on a somewhat limited investigation using Maple we have the following:
 \begin{conj}
 \label{conj:G(x) 3 primes}
 Suppose $x=p_1p_2p_3$ with $2 <p_1< p_2 < p_3$.
 Then
 $$
G(x)= \begin{cases} G(x-1) & \text{if  } p_1p_2>p_3, \\
G(x-1)+1 & \text{if  } p_1p_2<p_3.
\end{cases}
$$
 \end{conj}

We can also examine the cardinality of the set
 $$\cF(x):=\left\{n:\fl{\frac{x}{n}}\text{ is prime}\right\}.$$
 This might more naturally be thought of as the cardinality of the subsequence $(\cF_{n_k})$ created from the sequence $(\cF_n)_{n=1}^x, \, \cF_n=\fl{x/n}$, where you retain $n$ for which $\cF_n$ is prime and remove $n$ for which $\cF_n$ is not prime. For example, we have $$\cF(10)=\{2,3,4,5\},$$ whilst it is more natural to think of the sequence (for $x=10$) $$(\cF_{n_k})=5,3,2,2.$$ Of course, the cardinalities are the same.
 The cardinality of $\cF(x)$ (or of $(\cF_{n_k})$) can be obtained by substituting $f(m)= \textbf{1}_\mathbb{P}(m)$ into \eqref{eq:f sum} and using recent results from Wu \cite{Wu} or from  Zhai \cite{Zha}. As is usual,  $\textbf{1}_\mathbb{P}(m)=1$ if $m$ is prime and 0 otherwise.
We obtain the following:
 \begin{thm}
\label{thm:n prime}
Let $F(x):=|\cF(x)|$.Then
$$F(x)=\mathcal{P}x+O\(x^{1/2}\),$$
where
$$\mathcal{P}=\sum_{n=1}^\infty \frac{\textbf{1}_\mathbb{P}(n)}{n(n+1)}=\sum_p \frac1{p(p+1)}\cong 0.330230.$$
\end{thm}
We can use an alternate elementary approach, without exponential sums, to arrive at a result with a slightly better lower bound. Specifically
\begin{thm}
\label{thm:n prime elementary}
There exists calculable constants $A_1$ and $A_2$ such that for all $x$,
$$\mathcal{P}x -\frac{A_1\sqrt{x}}{\log x} \le F(x) \le \mathcal{P}x+A_2\sqrt{x}.$$
\end{thm}

The methodology of Theorem \ref{thm:n prime} can be utilised for all indicator functions since these functions are all bounded by 1. For example, we state, but do not prove, the following:
\begin{thm}
We have
$$\left\{n:\fl{\frac{x}{n}}\text{ is a prime power}\right\}=\mathcal{D}x +O\(x^{1/2}\),$$
where
$$\mathcal{D}=\sum_{n=p^k} \frac1{n(n+1)}\cong 0.41382.$$
\end{thm}

Throughout we use $p$, with or without subscript, to denote a prime number.
The notation $f(x) = O(g(x))$  or $f(x) \ll g(x)$ is
equivalent to the assertion that there exists a constant $c>0$ such that $|f(x)|\le c|g(x)|$ for all $x$.
As is normal, we denote by $\Lambda$ the von Mangoldt function.

\section{Proof of Theorem \ref{thm:G(x)}}
We have
$$G(x)=\left|\left\{p: p \le x, p=\fl{\frac{x}{n}}\text{ for some } 1 \le n \le x\right\} \right|.$$
If $\fl{\frac{x}{n}}=p$ then
$$\frac{x}{p+1} < n \le \frac{x}{p}$$ and such an $n$ will exist if
$\fl{\frac{x}{p}}-\fl{\frac{x}{p+1}}>0$. So
$$G(x)=\sum_{p \le x} \delta\(\fl{\frac{x}{p}} -\fl{\frac{x}{p+1}}>0\),$$
where $\delta=1$ if the statement is true and 0 otherwise.
Let
\begin{align}
\label{eq:G}
G(x)&=G_1(x)+G_2(x)+G_3(x)+G_4(x),
\end{align}
where
$$G_1= \sum _{p < b}, \,\, G_2=\sum_{b \le p \le \sqrt{x}},\,\, G_3=\sum_{\sqrt{x} < p \le x^{34/67}},\,\, G_4=\sum_{x^{34/67} < p \le x},$$
and
$$b=\frac{\sqrt{4x+1}-1}{2}=\sqrt{x}+O(1).$$

For $G_1(x)$ the condition is always satisfied, since for $p < b$ we have
$$\fl{\frac{x}{p}}-\fl{\frac{x}{p+1}}>\frac{x}{p}-\frac{x}{p+1}-1=\frac{x}{p(p+1)}-1>0.$$
So
\begin{align}
\label{eq:G1}
G_1(x)&=\sum_{p \le b}1=\pi(\sqrt{x})+O(1)=\frac{2\sqrt{x}}{\log x}+O\(\frac{\sqrt{x}}{(\log x)^2}\).
\end{align}

Trivially
\begin{align}
\label{eq:G2}
G_2(x)&=O(1).
\end{align}

Next, we estimate $G_4(x)$. If $p > x^{34/67}$ then $p=\fl{\frac{x}{n}}$ for some $n \le x^{33/67}.$ Since there can be at most $x^{33/67}$ values for $n$ we have
\begin{align}
\label{eq:G4}
G_4(x)&=O\(x^{33/67}\).
\end{align}
For $G_3(x)$ (and $G_4(x)$) $p$ is large enough that $\fl{\frac{x}{p}}-\fl{\frac{x}{p+1}}$ can only equal 0 or 1.
So
$$G_3(x)=\sum_{\sqrt{x} < p \le x^{34/67}} \delta\(\fl{\frac{x}{p}} -\fl{\frac{x}{p+1}}>0\)=\sum_{\sqrt{x} < p \le x^{34/67}} \(\fl{\frac{x}{p}} -\fl{\frac{x}{p+1}}\).$$
Then, using $\psi(x)=x-\fl{x}-\frac1{2}$,
\begin{align}
\label{eq:G3}
G_3(x)&=x \sum_{\sqrt{x}<p\le x^{34/67}}\frac1{p(p+1)}+\sum_{\sqrt{x}<p\le x^{34/67}}\(\psi\(\frac{x}{p+1}\)-\psi\(\frac{x}{p}\)\).
\end{align}
Using partial summation and the Prime Number Theorem we have, for the first sum,
\begin{align}
\label{eq:G3 first}
x \sum_{\sqrt{x}<p\le x^{34/67}}\frac1{p(p+1)}&=x\sum_{\sqrt{x}<n\le x^{34/67}}\frac{\textbf{1}_\mathbb{P}(n)}{p(p+1)}\notag\\
&=
\end{align}

For the second sum of $G_3(x)$ we will use the following (\cite[Theorem 6.25]{Bor3}):
\begin{lem}
\label{lem:Bor}
Let $\delta \in [0,1], x \ge 1$ be a large real number and $R$, $R_1$ be positive integers such that $1 \le R \le R_1 \le 2R \le x^{2/3}$. Then , for all $\epsilon \in (0,\frac1{2}]$
$$x^{-\epsilon}\sum_{R \le n \le R_1} \Lambda(n) \psi\(\frac{x}{n+\delta}\)\ll \(x^2R^{33}\)^{1/38}+\(x^2R^{19}\)^{1/24}\(x^3R^2\)^{1/9}+\(x^3R^{-1}\)^{1/6}+R^{5/6}.$$
\end{lem}

Returning to the second sum of $G_3(x)$ and using the Lemma we have
\begin{align*}
\sum_{\sqrt{x}<p\le x^{34/67}}\(\psi\(\frac{x}{p+1}\)-\psi\(\frac{x}{p}\)\)
&\le \left|\sum_{\sqrt{x}<p\le x^{34/67}}\psi\(\frac{x}{p+1}\)\right|+ \left|\sum_{\sqrt{x}<p\le x^{34/67}}\psi\(\frac{x}{p}\)\right|.
\end{align*}
We now bound the sum involving $\psi(\frac{x}{p})$. The calculations for the sum involving $\psi(\frac{x}{p+1})$ is virtually identical. Let $m \sim N$ denote the inequalities $N < m \le 2N$.
We have
\begin{align*}
\sum_{\sqrt{x}<p\le x^{34/67}}\psi\(\frac{x}{p}\)&\ll \max_{\sqrt{x}<N \le x^{34/67}}\left|\sum_{p \sim N} \psi\(\frac{x}{p}\)\right|\log x.
\end{align*}
Next, using Abel summation,
\begin{align*}
\left|\sum_{p \sim N} \psi\(\frac{x}{p}\)\right|&=\left|\sum_{p \sim N}\(\frac1{\log p}\times \psi\(\frac{x}{p}\) \log p \)\right|\\
&\le \frac{2}{\log N} \max_{N \le N_2 \le N_1}\left|\sum _{N <p \le N_2} \psi\(\frac{x}{p}\)\log p+\textbf{1}_p(N) \psi\(
\frac{x}{N}\log N\)\right|\\
&\le \frac{2}{\log N} \max_{N \le N_2 \le N_1}\left\{\left|\sum_{N<n \le N_2}\Lambda(n)\psi\(\frac{x}{n}\)\right|+|R(N)|+\log N\right\},
\end{align*}
where
$$|R(N)| \le \sum_{\sqrt{N} < p \le \sqrt{N_2}}\log p \sum_{2 \le a \le \frac{\log N_2}{\log p}} 1<2 \sqrt{N}.$$
Using  Lemma \ref{lem:Bor} with $N_1=N_2=x^{34/67}$ and $N=\sqrt{x}$ we obtain
\begin{align}
\label{eq:G3 second}
\sum_{\sqrt{x}<p\le x^{34/67}}\psi\(\frac{x}{p}\)&\ll x^{\frac{1256}{2546}+\epsilon}.
\end{align}
Substituting \eqref{eq:G3 first} and \eqref{eq:G3 second} into \eqref{eq:G3} and we see that
$$G_3(x)=\frac{2 \sqrt{x}}{\log x}+O\(\frac{\sqrt{x}}{(\log x)^2}\),$$
and substituting this equation and
 \eqref{eq:G1}, \eqref{eq:G2} and \eqref{eq:G4} into \eqref{eq:G} completes the proof.

\section{Proof of Theorem \ref{thm:adams}}
Let $x$ be any prime not equal to 3. Suppose $p \in \cG(x)$ but $p\ne x$. So for some $n$ we have $\fl{x/n}=p$. As $x$ is a prime we have $x=np+u$ where $1 \le u \le n-1$. So $x-1=np+u-1$ from which
$$\fl{\frac{x-1}{n}}=\fl{\frac{np+u-1}{n}}=\fl{p+\frac{u-1}{n}}=p.$$
Thus $p \in \cG(x-1)$.

Conversely, suppose $p \in \cG(x-1)$.  So $\fl{(x-1)/n}=p$ for some $n$. So $x-1=np+u$ where $0 \le u \le n-1$. But $u \ne n-1$, for then $x=np+n=n(p+1)$, which contradicts the supposition that $x$ is prime. Thus $x-1=np+u$ with $0 \le u \le n-2$, and then
$$\fl{\frac{x}{n}}=\fl{\frac{np+u+1}{n}}=\fl{p+\frac{u+1}{n}}=p.$$
Therefore $p \in \cG(x)$.

We conclude that there is one-to-one correspondence between an $p\ne x \in \cG(x)$ and $p \ne x \in \cG(x-1)$.
Noting that we have $x \in \cG(x)$ but $x \not \in \cG(x-1)$ concludes the proof.

\section{Proof of Theorem \ref{thm:G(x) semiprime}}
We have $x=pq$ with  $p,q$ odd primes, not necessarily distinct. Without loss of generality assume $p \le q$.

\textit{Case 1:}Suppose that $r \in \cG(x)$ with $r \ne p,q$.
So $x=nr+u$ with $0 \le u \le n-1$. But if $u=0$ then $x=nr$ which is impossible since $r \ne p,q$. So $1 \le u \le n-1$ and thus
$$\fl{\frac{x-1}{n}}=\fl{\frac{nr+u-1}{n}}=\fl{r+\frac{u-1}{n}}=r.$$
So $r \in \cG(x)$.
Conversely, suppose $r \in \cG(x-1)$. So $x-1=nr+u$ with $0 \le u \le n-1$. But if $u=n-1$ then $x=nr+n-1+1=(n+1)r$ which is again impossible since $r \ne p,q$. So $0 \le u \le n-2$ and then
$$\fl{\frac{x}{n}}=\fl{\frac{nr+u+1}{n}}=\fl{r+\frac{u+1}{n}}=r.$$
So $r \in \cG(x-1)$.

\textit{Case 2:} Suppose that $r \in \cG(x)$ with $r=p=q$.
Since $r=\fl{x/r}$ we have $r \in \cG(x)$. But
$$\fl{\frac{x-1}{r}}=\fl{r-\frac1{r}}=r-1$$
and
$$\fl{\frac{x-1}{r-1}}=\fl{\frac{r^2-1}{r-1}}=r+1.$$
So $r \not\in \cG(x-1).$

\textit{Case 3:} Suppose that $r \in \cG(x)$ with $r=p \ne q$ or $r=q \ne p$.
Recall $p \le  q$. It is clear that $p \in \cG(x)$ and $q \in \cG(x)$.
Then
$$\fl{\frac{x-1}{q-1}}=\fl{\frac{pq-1}{q-1}}=\fl{p+\frac{p-1}{q-1}}=p,$$
so $p \in \cG(x-1)$.
But
$$\fl{\frac{x-1}{p}}=\fl{\frac{pq-1}{p}}=\fl{q-\frac{1}{p}}=q-1,$$
and
$$\fl{\frac{x-1}{p-1}}=\fl{\frac{pq-1}{p-1}}=\fl{q+\frac{q-1}{p-1}}>q.$$
So $q \not \in \cG(x-1)$.

Reviewing the three cases we see that $G(x)=G(x-1)+1$, which proves the theorem.

\section{Proof of Theorem \ref{thm:n prime}}
We have
$$F(x)=\sum_{n \le x} \textbf{1}_\mathbb{P}\(\fl{\frac{x}{n}}\).$$
We will require the following, proven independently by Wu \cite[Theorem]{Wu} and Zhai \cite[Theorem 1]{Zha}:
\begin{lemma}
\label{thm:Sfx}
Let $f$ be a complex-valued arithmetic function such that $f(n) \ll n^{\alpha}(\log n)^{\theta}$ for some $\alpha \in [0,1)$ and $\theta \ge 0$.  Then
$$\sum_{n \leqslant x} f \(\fl{x/n} \) = x \sum_{n=1}^{\infty} \frac{f(n)}{n(n+1)} + O \( x^{\frac{1}{2}(\alpha+1)}(\log x)^{\theta}\).$$
\end{lemma}
Using this lemma with $\alpha=0$ and  $\theta=0$ we have
$$F(x)=\mathcal{P}x+O\(x^{1/2}\),$$
where
$$\mathcal{P}=\sum_{n=1}^\infty \frac{\textbf{1}_\mathbb{P}(n)}{n(n+1)}=\sum_p \frac1{p(p+1)}\cong 0.330230,$$
completing the proof.

We note, in passing, that
$$\sum_p \frac1{p(p+1)}=\sum_{s=2}^\infty \sum_p \frac{(-1)^s}{p^s}=\sum_{s=2}^\infty\sum_{n=1}^\infty \mu(n) \frac{\log \zeta(ns)}{n}.$$

\section{Proof of Theorem \ref{thm:n prime elementary}}
Let $\mathbb{P}$ be the set of (positive) primes and $\overline{\mathbb{P}}$ be the set of (positive) non-primes. We create upper and lower bounds for $F(x)$ from the set
$$\cC(x):=\{n: 1 \le n \le x\}.$$
The set $\cF(x)$ is simply the set $\cC(x)$ after removing all $n$ such that $\fl{x/n}$ is a non-prime from $\cC(x)$.
For the upper bound we truncate the process by removing from $\cC(x)$ only those $n$ such that $\fl{x/n}$ is a non-prime less than or equal to $\sqrt{x}$.
In total we remove
$$\sum_{\substack{c \in \overline{\mathbb{P}}\\c \le \sqrt{x}}}\(\fl{\frac{x}{c}}-\fl{\frac{x}{c+1}}\)>\sum_{\substack{c \in \overline{\mathbb{P}}\\c \le \sqrt{x}}}\(\frac{x}{c(c+1)}-1\)$$
values of $n$. So
\begin{align*}
F(x)&< x-\sum_{\substack{c \in \overline{\mathbb{P}}\\c \le \sqrt{x}}}\(\frac{x}{c(c+1)}-1\)\\
&= x-\sum_{\substack{c \in \overline{\mathbb{P}}\\c \le \sqrt{x}}}\frac{x}{c(c+1)}+\sum_{\substack{c \in \overline{\mathbb{P}}\\c \le \sqrt{x}}}1\\
&=x-x\(\sum_{c \le \sqrt{x}} \frac1{c(c+1)}-\sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}}\frac{1}{c(c+1)}\)+\sqrt{x}-\pi(\sqrt{x})\\
&=x-\frac{x\sqrt{x}}{\sqrt{x}+1} +x\sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}}\frac{1}{c(c+1)}+O\(\sqrt{x}\)\\
&=\frac{x}{\sqrt{x}+1}+x\sum_{p}\frac1{p(p+1)} -x\sum_{p>\sqrt{x}}\frac1{p(p+1)} +O\(\sqrt{x}\).
\end{align*}
Then
$$\sum_{p>\sqrt{x}}\frac1{p(p+1)}\le \sum_{n>\sqrt{x}}\frac1{n(n+1)}=O\(\frac1{\sqrt{x}}\),$$
and so
\begin{align}
F(x)& \le \mathcal{P}x+O\(\sqrt{x}\).
\end{align}
For the lower bound we add up the number of $n$ for which $\fl{x/n}$ is a prime less than or equal to $\sqrt{x}$.
Then, using $\pi(m)=\frac{m}{\log m} + O\(\frac{m}{(\log m)^2}\)$ and Riemann-Stieltjes integration,
\begin{align*}
F(x)&=\sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}}\(\fl{\frac{x}{c}}-\fl{\frac{x}{c+1}}\)
\ge \sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}} \(\frac{x}{c(c+1)}+1\)\\
&=x\sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}} \frac{1}{c(c+1)}+\sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}}1
=x\(\mathcal{P}-\sum_{p > \sqrt{x}}\frac{1}{p(p+1)}\)+\sum_{\substack{c \in \mathbb{P}\\c \le \sqrt{x}}}1\\
&=\mathcal{P}x-x \sum_{p>\sqrt{x}}\frac1{p(p+1)}+O\(\frac{\sqrt{x}}{\log x}\)=\mathcal{P}x+O\(\frac{\sqrt{x}}{\log x}\),
\end{align*}
completing the proof.

\section{Acknowledgements}
I thank Joshua Stucky who sketched out the proof of the asymptotic formula for $G(x)$. I thank Olivier Bordell\`es for assisting in the proof of Theorem \ref{thm:G(x)}. I thank Igor Shparlinski for some initial discussions. I thank Stan Wagon for pointing out that some of the inequalities in Conjecture 4 should be changed.

\makeatletter
\renewcommand{\@biblabel}[1]{[#1]\hfill}
\makeatother

\end{document}